Kobe and Shaq leave their home traveling in opposite directions on a straight road. Kobe drives 20 miles per hour faster that Shaq. After 4 hours they are 250 miles apart. Find their rates.
Answer by richwmiller(9144)

(20+x)*4+x*4=250;
80+8x=250
8x=170
x=170/8
x=85/4